OpenStudy (jojo12g):

A bag contains 5 black marbles and 4 white marbles. A spinner has 3 equal sections labeled 1 through 3. What is the probability of drawing a black marble from the bag and spinning a 1 on the spinner? A.5/9 B.1/2 C.2/9 D.5/27

OpenStudy (anonymous):

Total number of marbles = 9 (5/9) are black (4/9) are white Prob. of spinning a 1 = (1/3) (5/9)*(1/3) = 5/27

OpenStudy (jojo12g):

thanks

OpenStudy (anonymous):

np:)